OpenStudy (anonymous):

What are the diameters of the circles with areas 1) 201.1 2) 78.5

geerky42 (geerky42):

\(A = \pi r^2\) So since radius is half the diameter, then we have \(A = \pi\left(\dfrac{d}{2}\right)^2\), right? plug in area then try to isolate d.